the length of a rectangle is 5cm less than twice the width. The area is 75 square cm. Find the length and the width

Answer by fractalier(6550)

Area A = LW = 75 cm^2. We also have
L = 2W - 5
Substituting this we have
(2W - 5)W = 75
2W^2 - 5W - 75 = 0
(2W - 15)(W + 5) = 0
so that
2W = 15
W = 7.5 cm
and
L = 2(7.5) - 5 = 10 cm